OpenStudy (akohl):

A tennis ball is released from a height of 4.0 m above the floor. After its third bounce off the floor, it reaches a height of 183 cm. What percentage of its mechanical energy does it lose with each bounce? 15% 27% 20% 23%

OpenStudy (anonymous):

E = mgh the percentage = E1/E2= h/h=1.83/4=0.4575 and it make 3 to reach that height, so divide it by 3 = 0.15= 15%.

OpenStudy (akohl):

Thanks

OpenStudy (anonymous):

you are welcome.